# Effective grand-canonical description of condensation in negative-temperature regimes

### v1

@@ Abstract
The observation of negative-temperature states in the localized phase of the the Discrete Nonlinear Schrödinger (DNLS) equation has challenged statistical mechanics for a long time. For isolated systems, they can emerge as stationary extended states through a large-deviation mechanism occurring for finite sizes, while they are formally unstable in grand-canonical setups, being associated to an unlimited growth of the condensed fraction. Here, we show that negative-temperature states in open setups are metastable but their lifetime $\tau$ is exponentially long with the temperature, $\tau\approx\exp(\lambda|T|)$ (for $T<0$ and $\lambda>0$). More precisely, we find that condensation on a given site (i.e. emergence of a tall discrete breather) can advance only once a critical mass has been accumulated therein. This result has been obtained by combining the development of an effective grand-canonical formalism with the implementation of suitable heat baths. A general expression for $\lambda$ is obtained in the case of a simplified stochastic model of non-interacting particles. In the DNLS model, the presence of an adiabatic invariant, makes $\lambda$ even larger because of the resulting freezing of the breather dynamics. This mechanism, based on the existence of two conservation laws, provides a new perspective over the statistical description of condensation processes.
@@
Energy limitation can also arise indirectly in the presence of additional conservation laws. The discrete nonlinear Schrödinger (DNLS) equation, the subject of this Letter, is perhaps the most relevant such model Kevrekidis (2009). It is used to study many physical phenomena involving propagation of nonlinear waves in discrete media, from ultracold gases Trombettoni and Smerzi (2001); Livi et al. (2006); Hennig et al. (2010) in optical lattices to arrays of optical waveguides Christodoulides and Joseph (1988); Lederer et al. (2008). In the DNLS model, energy and norm (or mass) are conserved. For any given mass density, one can construct states of arbitrary energy, but above a critical energy density (corresponding to a so-called infinite temperature), the typical states involve the “condensation” of a finite fraction of energy on a few sites. In other words, the entropy decreases upon increasing energy: the standard signature of a NT.
The interplay between nonlinear dynamics and condensation makes the statistical mechanical description of macroscopic NT states particularly challenging. In spite of the compelling theoretical arguments predicting localization, DNLS microcanonical simulations performed above the infinite-temperature line have shown the emergence of “stationary” spatially homogeneous states Iubini et al. (2013a). Moreover, dynamical studies based on excursion-time distributions Mithun et al. (2018) revealed the occurrence of a transition from ergodic to nonergodic (localized) behavior, strictly above the infinite-temperature line. In the latter regime, localized fast rotations known as discrete breathers induce extremely slow relaxation phenomena Ng et al. (2009); Eckmann and Wayne (2018); Iubini et al. (2019). A partial solution of this paradox was offered in Gradenigo et al. (2021a, b), where it was proven that finite-size corrections are so strong that condensation is inhibited in a strip above the infinite-temperature line. The strip width decreases with the chain length $N$ as $N^{-1/3}$, but is large enough to justify some numerical observations of stationary “non-Gibbsian” states.
@@
So far, this scenario has been discussed in isolated systems; the regime in which mass and energy are exchanged with NT reservoirs received much less attention Baldovin et al. (2021). In this Letter, we revisit the problem by introducing an effective grand-canonical measure, which replaces the non-normalizable standard definition Rasmussen et al. (2000): as a result, we identify a novel class of metastable, delocalized NT states. In particular, we find that the condensation of energy on a single site (breather growth) is dynamically enabled only above a mass threshold $c_{max}\sim|T|$. A properly truncated distribution adequately describes the metastable regime and predicts exponentially long (with $|T|$) lifetimes. The predictions are confirmed by numerical simulations performed with suitable heat baths (after including some precautions to get rid of non-normalizable probabilities Aghion et al. (2019, 2020)). This mechanism, strengthened by the dynamical slowing down of breather states Iubini et al. (2019), clarifies that condensation instabilities can be de facto eluded, making NT states physically accessible also in open setups Kurnosov et al. (2024); Baudin et al. (2023); Iubini et al. (2017a).
@@
where $z_{n}$ is a complex variable, $n=1,\dots,N$ is the index of the lattice site and suitable boundary conditions are assumed (see below). The model has two conserved quantities: the total mass $A=\sum_{n}|z_{n}|^{2}$ and the total energy
@@
The equilibrium behavior of the model can be described in terms of the mass density $a=A/N$ and the energy density $h=H/N$. Equivalently, the inverse temperature $\beta$ and chemical potential $\mu$ Rasmussen et al. (2000) can be used within the “grand-canonical” representation. Below the infinite-temperature line $h_{c}=2a^{2}$, homogeneous positive-temperature states are found, well described by the Gibbs probability distribution $P_{G}(z_{n})\sim\exp(-\beta H+\beta\mu A)$. Above this line (i.e. for $\beta<0$), the grand-canonical picture is formally ill-defined due to the divergence of $P_{G}(z_{n})$ for arbitrarily large $|z_{n}|$.
@@
Close to $\beta=0$, the hopping energy $H_{int}=\sum_{n}(z_{n}z_{n+1}^{*}+z_{n}^{*}z_{n+1})$ can be neglected in Eq. (2Arezzo et al. (2022). The resulting model, called “C2C” Iubini et al. (2014); Szavits-Nossan et al. (2014); Gradenigo et al. (2021a); Gotti et al. (2021); Arezzo et al. (2022); Gotti et al. (2022), displays a phase-diagram very similar to that of the DNLS equation. It can be implemented as a stochastic evolution involving real variables $c_{n}=|z_{n}|^{2}\geq 0$. A triple of consecutive sites is randomly selected, and the corresponding variables $(c_{n-1},c_{n},c_{n+1})$ are updated in such a way that their sum and the sum of their squares are both conserved. The admissible configurations lie along a circle and the selection is randomly made by assuming a uniform angle distribution. Depending on the initial configuration, the “legal” states (with positive masses) may be restricted to three distinct arcs: in this case, the new state is selected within the same arc of the initial one. Detailed balance is satisfied.
@@
The average densities $a=\langle A\rangle/N$ and $h=\langle H\rangle/N$ can be thereby derived from the standard grand-canonical relations, $a=\partial_{m}\log(Z_{r})$, $h=\partial_{T}\log(Z_{r})/T^{2}$, obtaining sup
$\begin{cases}a&=\sqrt{|T|}\tilde{D}^{-1}\left(m\sqrt{|T|}/2\right)/2-m|T|/2\\ h&=-(ma+1)|T|/2\,.\end{cases}$ (5)
For $|T|\gg 1$, these expressions reduce to
@@
which hold independently of the sign of $\beta$. Hence, as long as $c<c_{max}$, the stationary properties change smoothly while crossing the infinite-temperature line. This result is valid also in the thermodynamic limit, but is limited to finite observation times which depend on the (size-dependent) escape time from the interval $[0,c_{max}]$. We now discuss the metastability of this regime.
“Stationary” dynamics
If we formally view $P(c)$ as the outcome of a Langevin process, we can interpret $U(c)\simeq-\ln P(c)=\beta c^{2}-mc$ as an effective potential (complemented by the condition $U(0)=+\infty$). A few samples are presented in Fig. 1: $U(c)$ is binding for $\beta>0$ (see the black curve). For $\beta=0$, $U(c)$ is still binding provided that $m<0$: see the green straight curve, which corresponds to a pure exponential distribution. For $\beta<0$, $U(c)$ is no longer binding, but it is confining for $c<c_{max}$ (the red curve in Fig. 1 corresponds to $\beta=-0.1$).
@@
So far, the interpretation of $U(c)$ is formal. In order to validate its meaning, we consider the restriction of the C2C model to a single triplet $(c_{-},c,c_{+})$, where $c_{-}$ and $c_{+}$ are independent random variables, both distributed according to Eq. (3). It is straightforward to conclude that also the distribution of $c$ is described by Eq. (3) for the same $\beta$ and $m$ values. For NT, $P(c)$ is formally unbounded. However, consistently with our effective grand-canonical approach, it makes sense, slightly above the $\beta=0$ line, to cut the range of allowed $c$ values at $c_{max}$ (i.e. at a minimum of $P(c)$).
Next, we approximate this stochastic system with a nonlinear Langevin process,
$\dot{c}=F(c)+\sqrt{2D(c)}\xi\;,$ (7)
where the effective deterministic drift $F(c)$ and the diffusion coefficient $D(c)$ ($\xi$ is a white noise with unit variance) are empirically defined from numerical simulations
Here $c^{\prime}$ is the iterate of $c$ as obtained by implementing the C2C model, while the angular brackets denote an average over different realizations of $c_{-}$, $c_{+}$ and of the angle selected in the implementation of the C2C rule.
The results for infinite temperature and $T=-100$ are plotted in Fig. 2. In both cases $F$ is negative for small $c$ values, indicating that small breathers are stable. A better description is obtained by looking at $F/D$ Gardiner (1985); first of all, we can appreciate that for $\beta=0$, the drift $F(c)$ is negative for all $c$ values (see panel (a)) in agreement with the expectation that breathers of any amplitude are eventually absorbed. For $T=-100$, the drift crosses zero at $c=50$, becoming positive (see the green line in pnale (b)) in agreement with the theoretical expectation $c_{max}=mT/2$ ($m=-1$). On a yet more quantitative level, the asymptotic linear growth of $F/D$ matches the linear dependence of $-U^{\prime}(c)=2\beta c-m$, thus showing that the Langevin equation reproduces the behavior of the C2C model for large $c$ values. In particular, only breathers with $c>c_{max}$ tend to explode, or, equivalently, represent seeds of condensation.
Now, we focus on the lifetime of the metastable states, starting from their stationarity. In Fig. 3 we present the distribution of $c$ values obtained by averaging over 1000 different trajectories until they reach the maximum of the potential at $c_{max}$ (the initial condition having been fixed in $c=1$).
@@
Thanks to the correspondence with the Langevin model, the lifetimes of the metastable state can be estimated via the Kramers formula Risken (1989) for the escape through a barrier located in $c_{max}$, $\tau\approx\exp(\Delta U)/\sqrt{D_{0}D_{1}}$, where $\Delta U$ is the barrier height, while $D_{0}\equiv D(0)$ and $D_{1}\equiv D(c_{max})$ are the diffusion coefficients in the bottom of the valley and at the saddle, respectively. Since $\Delta U=m^{2}|T|/4$, $\tau\approx\exp(\lambda|T|)$, with $\lambda=m^{2}/4$. Numerical simulations performed for $m=-1$ and $T\in[-50,-25]$ yield $\lambda\approx 0.365$ to be compared with the theoretical prediction 0.25. Large part of the discrepancy can be attributed to the indirect dependence of $D_{1}$ on $T$. In fact, as visible in both panels of Fig. 2, $D$ decreases with $c$, while $c_{max}=-T/2$. The effective dependence follows a power law Iubini et al. (2014, 2017b) and thus it does not affect the asymptotic scaling behavior; however, in the range of temperatures here explored, $D_{1}$ contributes an increase of $\lambda$ by about 0.07, bringing the theoretical prediction closer to the observed value (0.32 vs 0.365). The remaining discrepancy is presumably due to the imperfect description of the stochastic dynamics in terms of a continuous nonlinear Langevin process.
@@
Given the above mentioned continuity with the high-temperature region, it makes sense to extend the Langevin scheme proposed in Iubini et al. (2013b) to model the interaction of a DNLS chain with NT reservoirs
$\dot{z}_{n}\!=\!(\gamma_{n}-i)(-2|z_{n}|^{2}z_{n}\!-\!z_{n+1}\!-\!z_{n-1})+\gamma_{n}\mu z_{n}+\sqrt{\gamma_{n}T}\xi_{n}\,$ (9)
where $\gamma_{n}$ is the coupling strength and $\xi_{n}$ a complex Gaussian noise with zero mean and unit variance. While $\gamma_{n}$ is positive for $T>0$, it changes sign at NT. This property is in accord with the general symmetries expected for NT reservoirs Baldovin et al. (2021); $\gamma_{n}<0$ indicates that the sign of nonconservative forces is reversed with respect to the $\beta>0$ case. If we neglect nearest-neighbor interactions in Eq. (9), the single-site non-conservative force $\tilde{F}(z)\simeq\gamma_{n}z_{n}\left(\mu-2|z_{n}|^{2}\right)$ is destabilizing for $|z_{n}|^{2}\geq c_{max}$, as predicted by the single-particle probability $P(c)$, Eq. (3). Similar results had been obtained while studying the stability of breathers in interaction with a bath of linear waves Rumpf (2007, 2009).
In the following we assume periodic boundary conditions ($z_{N+1}=z_{1}$) and thermalize a single site, $\gamma_{n}=\gamma\delta_{n,1}$. The inset of Fig. 4 shows the averages of $a$ and $h$ in a thermostatted DNLS chain with 20 lattice sites, for $T=-100$, $m=-1$ and $\gamma=-0.02$ (in this regime $\tau>10^{9}$ is much longer than the integration time). They are in good quantitative agreement with the truncated grand-canonical prediction Eq. (5) (dashed lines), as also confirmed by the direct extraction of mass distributions, see the upper inset of Fig. 3. This is not a priori obvious, since the theory neglects $H_{int}$. In the metastable $\beta<0$ region of interest for this work $H_{int}$ is indeed negligible sup .
@@
We now study the evolution of discrete breathers determining their drift and diffusion properties as a function of their amplitude $|z^{2}|$. The drift $F(c)$ is estimated by monitoring the evolution of a breather of initial mass $|z|^{2}=c$ in a short ($N=6$) chain thermostatted at a given NT in the antipodal position with respect to the breather site $n_{0}$. To improve numerical accuracy Politi et al. (2022), we monitor the square root of breather energy $\tilde{c}(t)=\{\,|z_{n_{0}}|^{4}+[z_{n_{0}}^{*}(z_{n_{0}+1}+z_{n_{0}-1})+c.c.]\,\}^{1/2}$, (dimensionally equivalent to the local mass), and compute $F(c)=\langle d\tilde{c}/dt\rangle$. Analogously, $D(c)=Var[\,\tilde{c}(t)\,]/2t$, where $Var(\cdot)$ denotes the variance of the signal. The red crosses reported in Fig. 4 show that the drift is negative for small $c$-values, indicating that such breathers are stable, consistently with the numerical evidence of pseudo stationary regimes. However, at variance with the stochastic C2C model, here $F(c)$ becomes indistinguishable from zero above $c\approx 17$. This slowing down is a manifestation of the adiabatic invariant (AI) unveiled in Iubini et al. (2019); Politi et al. (2022) in the positive-temperature region.
In order to bypass this obstacle, we have also analyzed a stochastic version of the DNLS equation (SDNLS), where the AI is absent. It amounts to including the sporadic update of the phase of a randomly chosen $z_{n}$ so as to leave the local interaction energy unchanged sup ; Iubini (2019). The SDNLS exhibits the same stationary properties as the DNLS model, but is characterized by a substantially faster breather dynamics.
@@
Simulations of the SDNLS chain for different temperatures show that the drift is now significantly different from 0 and quantitatively measurable (see circles and triangles in Fig. 4(c)). In particular, one can identify a crossing point above which the drift is positive. Also the diffusion coefficient can be estimated; the resulting numerical values of $F/D$ are presented in Fig. 4(d). They are relatively close to the theoretical predictions (see the dashed lines); the main difference is a rightward shift which implies that the numerically estimated saddle point $\tilde{c}_{max}$ is slightly larger than the expected value (on the basis of the C2C model). Anyway, the deviation $\tilde{c}_{max}-c_{max}$ does not increase upon increasing $|T|$ sup .
A fully quantitative analysis of the deterministic DNLS equation is not doable. However, we have made our best efforts to determine $F/D$ in a single point in the presence of an AI. We have performed a single long numerical simulation for $T=-34$ and an initial breather mass $c=21>c_{max}=17$ $N=12$: the evolution of $\tilde{c}(t)$ is reported in the inset of Fig. 4(d). It reveals large bursts and less visible but relevant blobs which all hinder a quantitative analysis, since these localized instabilities do not typically contribute to the overall growth and diffusion  [^1] They signal the emergence of “dimer” bound states when the energy temporarily spreads over two adjacent sites and yet strongly affect numerical estimates. After a lengthy manual removal of the most relevant bursts, we have obtained a cleaner curve  sup , whose analysis leads to $F/D\approx 0.38$, in semi-quantitative agreement with the theoretical expectation (see the uppermost red cross in Fig. 4(d)).
Finally, we consider the escape time out of the DNLS metastable regime. The major difference with the C2C model is the dependence of the diffusion coefficient $D$ on the mass contained in the breather. As shown in Iubini et al. (2019); Politi et al. (2022), the AI induces an exponential decrease, $D\approx\exp(-\alpha c)$. Hence, $D_{1}\simeq\exp(-\alpha mT/2)$, which, inserted into the Kramers formula, enhances the exponential growth rate of the lifetime, to $\lambda\approx m^{2}/4+\alpha|m|/4$. This implies (for $m=-1$), $\lambda\approx 0.5$. Direct numerical simulations performed for $N=7$ yield the results reported in Fig. 3, an exponential fit gives a rate 0.53 in good agreement with the theoretical expectation.

### v2

@@ Abstract
The observation of negative-temperature states in the localized phase of the Discrete Nonlinear Schrödinger (DNLS) equation has challenged statistical mechanics for a long time. For isolated systems, they can emerge as stationary extended states through a large-deviation mechanism occurring for finite sizes, while they are formally unstable in grand-canonical setups, being associated to an unlimited growth of the condensed fraction. Here, we show that negative-temperature states in open setups are metastable and their lifetime $\tau$ is exponentially long with the temperature, $\tau\approx\exp(\lambda|T|)$ (for $T<0$). A general expression for $\lambda$ is obtained in the case of a simplified stochastic model of non-interacting particles. In the DNLS model, the presence of an adiabatic invariant, makes $\lambda$ even larger because of the resulting freezing of the breather dynamics. This mechanism, based on the existence of two conservation laws, provides a new perspective over the statistical description of condensation processes.
@@
Energy limitation can also arise indirectly in the presence of additional conservation laws. The discrete nonlinear Schrödinger (DNLS) equation, the subject of this Letter, is perhaps the most relevant such model Kevrekidis (2009). It is used to study many physical phenomena involving propagation of nonlinear waves in discrete media, from ultracold gases Trombettoni and Smerzi (2001); Livi et al. (2006); Hennig et al. (2010) in optical lattices to arrays of optical waveguides Christodoulides and Joseph (1988); Lederer et al. (2008). In the DNLS model, energy and norm (or mass) are conserved. For any mass density, one can construct states of arbitrary energy, but above a critical energy density (corresponding to a so-called infinite temperature), the entropy decreases upon increasing energy: the standard signature of a NT. As a result, typical states involve the “condensation” of a finite fraction of energy on a few sites.
Nevertheless, in Gradenigo et al. (2021a, b), it was proven that finite-size corrections are so strong that condensation is inhibited in a strip of width $N^{-1/3}$ ($N$ being the chain length), above the infinite-temperature line. The strip is large enough to justify some numerical results, such as the observation of “stationary” spatially homogeneous states Iubini et al. (2013a) and of “non-Gibbsian” states Mithun et al. (2018) (the latter ones detected via a careful analysis of excursion-time distributions).
The complexity of the underlying regime can be traced back to the emergence of localized fast rotations known as discrete breathers, around which a sensible amount of energy can condense. In the past, breathers have been studied as single dynamical objects, by investigating the stability of energy clumps Trombettoni and Smerzi (2001); Hennig et al. (2016); Franzosi et al. (2011). Here, we aim at characterizing breather dynamics in a statistical-mechanics setup, by singling out the role of thermodynamic observables.
A statistical description of localization stability has been already produced for breathers interacting with a bath of waves in terms of a matching condition between breather frequency and the chemical potential of the wave background Rumpf (2007, 2009); Chen and Rumpf (2021). Furthermore, tall (i.e. rapidly rotating) breathers have been found to undergo extremely slow relaxation phenomena Ng et al. (2009); Eckmann and Wayne (2018); Iubini et al. (2019).
@@
So far, this problem was discussed in isolated systems. In this Letter, we study a setup where mass and energy are exchanged with NT reservoirs, by introducing an effective grand-canonical measure, which replaces the non-normalizable standard definition Rasmussen et al. (2000). As a result, we identify a novel class of extended NT states and describe the stationary regimes in terms of a properly truncated distribution. With the help of an effective Langevin equation (after including some precautions to get rid of non-normalizable probabilities Aghion et al. (2019, 2020)), the steady description is inserted into a dynamical context, which leads to two predictions: (i) condensation progresses (on average) only above a critical mass threshold $c_{max}\sim|T|$, in agreement with the results in Rumpf (2007, 2009); Chen and Rumpf (2021); (ii) the lifetime of the associated metastable state is exponentially long (with $|T|$). The lifetime is lenghtened by the dynamical slowing down of tall breathers Iubini et al. (2019), de facto eluding the condensation instabilities and making NT states physically accessible also in open setups Kurnosov et al. (2024); Baudin et al. (2023); Iubini et al. (2017a).
@@
where $z_{n}$ is a complex variable, $n=1,\dots,N$ is the index of the lattice site and suitable boundary conditions are assumed (see below). The model has two conserved quantities: the total mass $A=\sum_{n}|z_{n}|^{2}$ and the total energy  [^1] Sometime in the literature, the quartic term is preceded by a coefficient $\nu$ quantifying the nonlinearity Rasmussen et al. (2000). It can be eliminated via a suitable rescaling. Its explicit presence implies that the expression of the critical threshold $c_{max}$ should be divided by $\nu$ (see below).
@@
The equilibrium behavior of the model can be described in terms of the mass density $a=A/N$ and the energy density $h=H/N$. Equivalently, the inverse temperature $\beta$ and chemical potential $\mu$ Rasmussen et al. (2000) can be used within the “grand-canonical” representation. Below the infinite-temperature line $h_{c}=2a^{2}$, homogeneous positive-temperature states exista and are well described by the Gibbs probability distribution $P_{G}(z_{n})\sim\exp(-\beta H+\beta\mu A)$. Above this line (i.e. for $\beta<0$), the grand-canonical picture is formally ill-defined due to the divergence of $P_{G}(z_{n})$ for arbitrarily large $|z_{n}|$.
@@
Close to $\beta=0$, the hopping energy $H_{int}=\sum_{n}(z_{n}z_{n+1}^{*}+z_{n}^{*}z_{n+1})$ can be neglected in Eq. (2Arezzo et al. (2022). The resulting model, called “C2C” Iubini et al. (2014); Szavits-Nossan et al. (2014a); Gradenigo et al. (2021a); Gotti et al. (2021); Arezzo et al. (2022); Gotti et al. (2022), displays a phase-diagram very similar to that of the DNLS equation. It can be implemented as a stochastic evolution involving real variables $c_{n}=|z_{n}|^{2}\geq 0$ in such a way that their sum and the sum of their squares are both conserved sup .
@@
The average densities $a=\langle A\rangle/N$ and $h=\langle H\rangle/N$ can be derived from the grand-canonical relations, $a=\partial_{m}\log(Z_{r})$, $h=\partial_{T}\log(Z_{r})/T^{2}$ (see sup ). For $|T|\gg 1$, the expressions reduce to
@@
which hold independently of the sign of $\beta$. Hence, as long as $c<c_{max}$, the stationary properties (including those of spatial profiles) change smoothly while crossing the infinite-temperature line.
One can formally interpret $U(c)\simeq-\ln P(c)=\beta c^{2}-mc$ as the effective potential (complemented by the condition $U(0)=+\infty$) in an underlying Langevin process. A few instances are presented in Fig. 1: $U(c)$ is binding for $\beta>0$ (see the black curve). For $\beta=0$, $U(c)$ is still binding provided that $m<0$: see the green straight curve, which corresponds to a pure exponential distribution. For $\beta<0$, $U(c)$ is no longer binding, but it is confining for $c<c_{max}$ (the red curve in Fig. 1 corresponds to $\beta=-0.1$).
@@
In order to validate this formal interpretation, we consider the restriction of the C2C model to a single triplet $(c_{-},c,c_{+})$, where $c_{-}$ and $c_{+}$ are independent random variables, both distributed according to Eq. (3). It is straightforward to conclude that also the distribution of $c$ is described by Eq. (3) for the same $\beta$ and $m$ values. For NT, $P(c)$ is formally unbounded. However, consistently with our effective grand-canonical approach, it makes sense, slightly above the $\beta=0$ line, to cut the range of allowed $c$ values at $c_{max}$ (i.e. at a minimum of $P(c)$).
In Fig. 2 we present the distribution of $c$ values obtained by averaging over 1000 different trajectories sampled until they reach $c_{max}$ (the initial condition having been fixed in $c=1$).
@@
Effective Langevin description
Here, we approximate the model dynamics with the stochastic equation
$\dot{c}=F(c)+\sqrt{2D(c)}\xi\;,$ (6)
where the deterministic drift $F(c)$ and the diffusion coefficient $D(c)$ ($\xi$ is a white noise with unit variance) are empirically defined from numerical simulations
Here $c^{\prime}$ is the iterate of $c$ as obtained by implementing the C2C model. The angular brackets denote an average over all possible realizations of the thermal baths and implementations of the C2C rule.
In Fig. 3, we see that both for $\beta=0$ and $-0.01$, $F$ is negative for (not too) small $c$ values, indicating that small breathers are stable. $F/D$ Gardiner (1985) provides a better representation. For $\beta=0$, we see that the drift is still negative at large $c$ values (see panel (a)) in agreement with the expectation that breathers are eventually absorbed. For $\beta=-0.01$, the drift crosses zero at $c=50$, becoming positive (see the green line in panel (b)) in agreement with the theoretical expectation $c_{max}=mT/2$ ($m=-1$). On a yet more quantitative level, the asymptotic linear growth of $F/D$ matches the linear dependence of $-U^{\prime}(c)=2\beta c-m$, thus showing that the Langevin equation reproduces the behavior of the C2C model for large $c$ values. In particular, only breathers with $c>c_{max}$ tend to explode, or, equivalently, represent seeds of condensation.
Once $F$ and $D$ have been determined, one can estimate the lifetime of the metastable state via Kramers formula Risken (1989) for the escape through a barrier located in $c_{max}$, $\tau\approx\exp(\Delta U)/\sqrt{D_{0}D_{1}}$, where $\Delta U$ is the barrier height, while $D_{0}\equiv D(0)$ and $D_{1}\equiv D(c_{max})$ are the diffusion coefficients in the bottom of the valley and at the saddle, respectively. Since $\Delta U=m^{2}|T|/4$, $\tau\approx\exp(\lambda|T|)$, with $\lambda=m^{2}/4$. Numerical simulations performed for $m=-1$ and $T\in[-50,-25]$ yield $\lambda\approx 0.365$ to be compared with the theoretical prediction 0.25. Large part of the discrepancy can be attributed to the indirect dependence of $D_{1}$ on $T$. In fact, as visible in both panels of Fig. 3, $D$ decreases with $c$, while $c_{max}=-T/2$. The effective dependence follows a power law Iubini et al. (2014, 2017b) and thus it does not affect the asymptotic scaling behavior; however, in the range of temperatures here explored, $D_{1}$ contributes an increase of $\lambda$ by about 0.07, bringing the theoretical prediction closer to the observed value (0.32 vs 0.365). The remaining discrepancy is presumably due to the imperfect description of the stochastic dynamics in terms of a continuous nonlinear Langevin process.
@@
Given the above mentioned continuity with the high-temperature region, it makes sense to extend the scheme proposed in Iubini et al. (2013b) to model the interaction of a DNLS chain with NT reservoirs
$\dot{z}_{n}\!=\!(\gamma_{n}-i)(-2|z_{n}|^{2}z_{n}\!-\!z_{n+1}\!-\!z_{n-1})+\gamma_{n}\mu z_{n}+\sqrt{\gamma_{n}T}\xi\,$ (8)
where $\gamma_{n}=\delta_{n1}\gamma$ is the coupling strength and $\xi$ a complex Gaussian noise with zero mean and unit variance. Periodic boundary conditions ($z_{N+1}=z_{1}$) are assumed. While $\gamma$ is positive for $T>0$, it changes sign at NT. This property is in accord with the general symmetries expected for NT reservoirs Baldovin et al. (2021); $\gamma<0$ indicates that the sign of nonconservative forces is reversed with respect to the $\beta>0$ case. If we neglect the nearest-neighbor interactions of the thermostatted site, the evolution equation of the mass $c=|z_{1}|^{2}$ writes $\dot{c}=-2\gamma(2c^{2}-\mu)c+z^{*}\xi+z\xi^{*}$, a form similar to the Langevin Eq. (6), with the difference that this equation follows from a precise thermal-bath definition Iubini et al. (2013b) instead of being a mere conjecture. It is also interesting to notice that the effective force $\tilde{F}(c)=2\gamma c\left(\mu-2c^{2}\right)$ predicts a destabilization for $c^{2}\geq c_{max}$, consistent with the single-particle probability $P(c)$, Eq. (3).
Fig. 4 (a,b) shows the averages of $a$ and $h$ in a DNLS chain of length $N=20$, thermostatted for $T=-100$, $m=-1$ and $\gamma=-0.02$ (in this regime $\tau>10^{9}$ is much longer than the integration time). They are in good quantitative agreement with the truncated grand-canonical prediction Eq. (5) (dashed lines), as also confirmed by the direct extraction of mass distributions, see the upper inset of Fig. 2. This correspondence is not a priori obvious, since the theory neglects $H_{int}$. In the metastable $\beta<0$ region of interest for this work, $H_{int}$ is on average negligible sup .
@@
We now study the evolution of discrete breathers for different amplitudes $|z^{2}|$. At variance with the C2C model, here thermostats induce strong boundary effects. As previously shown in positive-temperature simulations Politi et al. (2022), the high frequencies naturally generated by the heat bath spuriously affect the DNLS dynamics (in particular the breather stability). Hence, Eq. (8) cannot account for the true dynamics of the background. Nevertheless, Eq. (6) provides again a good effective description. By determining the drift $F$ and diffusion $D$ of breathers of mass $|z|^{2}=c$ (located in the antipodal position with respect to the thermal bath) for different chain lengths, we have verified that a length $N=6$ suffices to kill unphysical high frequencies.
To improve numerical accuracy Politi et al. (2022), we monitor the square root of breather energy $\tilde{c}(t)=\{\,|z_{n_{0}}|^{4}+[z_{n_{0}}^{*}(z_{n_{0}+1}+z_{n_{0}-1})+c.c.]\,\}^{1/2}$, (dimensionally equivalent to the local mass), and compute $F(c)=\langle d\tilde{c}/dt\rangle$. Analogously, $D(c)=Var[\,\tilde{c}(t)\,]/2t$, where $Var(\cdot)$ denotes the variance of the signal. The red crosses reported in Fig. 4 show that the drift is negative for small $c$-values, indicating that such breathers are stable, consistently with the numerical evidence of pseudo stationary regimes. However, at variance with the stochastic C2C model, here $F(c)$ becomes indistinguishable from zero above $c\approx 17$. This slowing down is a manifestation of the adiabatic invariant unveiled in Iubini et al. (2019); Politi et al. (2022) in the positive-temperature region.
In order to bypass this obstacle, we have also analyzed a stochastic version of the DNLS equation (SDNLS), where the adiabatic invariant is absent. It amounts to including the sporadic update of the phase of a randomly chosen $z_{n}$ so as to leave the local interaction energy unchanged sup ; Iubini (2019). The SDNLS exhibits the same stationary properties as the DNLS model, but is characterized by a substantially faster breather dynamics.
@@
Simulations of the SDNLS chain show that the drift is now significantly different from 0 and quantitatively measurable (see circles and triangles in Fig. 4(c)). In particular, one can identify a crossing point above which the drift is positive. Also the diffusion coefficient can be estimated; the resulting numerical values of $F/D$ are presented in Fig. 4(d). They are relatively close to the theoretical predictions (see the dashed lines); the main difference is a rightward shift which implies that the numerically estimated saddle point $\tilde{c}_{max}$ is slightly larger than the expected value (on the basis of the C2C model). Anyway, the deviation $\tilde{c}_{max}-c_{max}$ does not increase upon increasing $|T|$ sup .
A fully quantitative analysis of the deterministic DNLS equation is not doable. However, we have made our best efforts to determine $F/D$ in a single point in the presence of an adiabatic invariant. We have performed a single long numerical simulation for $T=-34$ and an initial breather mass $c=21>c_{max}=17$ for $N=12$: the evolution of $\tilde{c}(t)$ is reported in the inset of Fig. 4(d). It reveals large bursts and less visible but relevant blobs which all hinder a quantitative analysis, since these localized instabilities do not typically contribute to the overall growth and diffusion  [^2] They signal the emergence of “dimer” bound states when the energy temporarily spreads over two adjacent sites and yet strongly affect numerical estimates. After a lengthy manual removal of the most relevant bursts, we have obtained a cleaner curve  sup , whose analysis leads to $F/D\approx 0.38$, in semi-quantitative agreement with the theoretical expectation (see the uppermost red cross in Fig. 4(d)).
Finally, we consider the escape time out of the DNLS metastable regime. The major difference with the C2C model is the dependence of the diffusion coefficient $D$ on the mass contained in the breather. As shown in Iubini et al. (2019); Politi et al. (2022), the adiabatic invariant induces an exponential decrease, $D\approx\exp(-\alpha c)$, with $\alpha\approx 1$. Hence, $D_{1}\simeq\exp(-\alpha mT/2)$, which, inserted into the Kramers formula, enhances the exponential growth rate of the lifetime, to $\lambda\approx m^{2}/4+\alpha|m|/4$. This implies (for $m=-1$), $\lambda\approx 0.5$. Direct numerical simulations performed for $N=7$ yield the results reported in Fig. 2; an exponential fit gives a rate 0.53 in good agreement with the theoretical expectation.